A pulley is used to pull a bale of hay 5.80m vertically up to try to get it into a hay loft. As the bale reaches the top, the rope breaks and the bale falls to the ground how fast was the bale falling


1
Expert's answer
2021-04-28T20:15:51-0400
"v^2=2gh\\v^2=2(9.8)(5.8)\\v=10.7\frac{m}{s}"